3At what temperature is the following reaction at e
Answer #1

Here we have the reaction

2Na_2O_2+2H_2O\rightarrow 4NaOH+O_2

And

\\ \Delta H^\circ =-109KJ=-109000J\\ \Delta S^\circ = -133.2J/K

At equilibrium free energy is zero that means \\ \Delta G^\circ =0 ,

As we know

\\ \Delta G^\circ =\Delta H^\circ -T\Delta S^\circ \\ \text{At equilibrium, we have }\\ \Delta G^\circ =0\\ \Delta H^\circ -T\Delta S^\circ=0\\ \Delta H^\circ=T\Delta S^\circ\\ T=\frac{\Delta H^\circ}{\Delta S^\circ}\\ T= \frac{-109KJ}{-133.2J/K}\\ T= \frac{-109\times 1000J}{-133.2J/K}\\ T= 818.31 K

Hence the temperature required at equilibrium is 818.31 K.